OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <6, -2>, v = <8, 24>

jimthompson5910 (jim_thompson5910):

Let u = <a,b> v = <c,d> the dot product of u and v is u dot v = a*c + b*d if the dot product of u and v is 0, then u is orthogonal (ie perpendicular) to v.

OpenStudy (anonymous):

it is orthogonal, correct?

jimthompson5910 (jim_thompson5910):

u is parallel to v if and only if u = k*v for some scalar k u = k*v <a,b> = k*<c,d> <a,b> = <k*c,k*d> a = k*c and b = k*d k = a/c and k = b/d a/c = b/d All that work above shows that IF a/c = b/d, then u and v are parallel. They either point in the same direction, or they point in complete opposite directions.

jimthompson5910 (jim_thompson5910):

`it is orthogonal, correct?` yes that is correct u dot v = a*c + b*d u dot v = 6*8 + (-2)*24 u dot v = 48 + (-48) u dot v = 0 so they are indeed orthogonal

OpenStudy (anonymous):

thank you!!!

jimthompson5910 (jim_thompson5910):

sure thing